OpenStudy (anonymous):

Solve by Factoring 1. 3x^2 + 7x - 40 = 0 2. 4x^2 -25 = 0

OpenStudy (anonymous):

1. 3x^2+15x-8x-40 = 0 or 3x(x+5) -8(x+5) = 0 or (3x-8)(x+5)=0 2. (2x)^2-5^2 =0 (2x+5)(2x-5) = 0